OpenStudy (anonymous):

Can someone help me solve 92 29/30 - 15 13/15? I know the common denominator would be 15...

OpenStudy (phi):

subtract the whole numbers subtract the fractions

ganeshie8 (ganeshie8):

\(\huge 92\frac{29}{30} - 15\frac{13}{15}\)

ganeshie8 (ganeshie8):

its like above ?

OpenStudy (anonymous):

yes, just like that!

ganeshie8 (ganeshie8):

you can write the same as : \(\huge (92 + \frac{29}{30}) - (15 + \frac{13}{15}) \)

ganeshie8 (ganeshie8):

you can group the whole numbers, and fractions separately

ganeshie8 (ganeshie8):

\(\huge (92 - 15) + (\frac{29}{30} - \frac{13}{15})\)

ganeshie8 (ganeshie8):

does this make sense so far ?

hero (hero):

You stole my method @ganeshie8, but I'm cool with it. :P

ganeshie8 (ganeshie8):

lol i stole your many methods :P

OpenStudy (anonymous):

okay so the fraction would be 16/15?

ganeshie8 (ganeshie8):

nopes... while subtracting fractions, you need to make denominators common

ganeshie8 (ganeshie8):

the fraction part : \(\huge (\frac {29}{30} - \frac{13}{15})\) \(\huge (\frac {29}{30} - \frac{13*2}{15*2})\) \(\huge (\frac {29}{30} - \frac{26}{30})\)

ganeshie8 (ganeshie8):

now since the denominators are same, you can subtract numerators !!

OpenStudy (anonymous):

oh okay, so the answer is 77 3/1? I'm confused with the 30

ganeshie8 (ganeshie8):

okay.... 30 is just the LCM or common-denominator 77 3/30 you can simplify further, 3 goes in 30

OpenStudy (anonymous):

77 1/10?

ganeshie8 (ganeshie8):

thats right ! good work :P btw, did you get that 30 thingy in fraction part... ?

OpenStudy (anonymous):

Oh I meant subtracting 30 - 30

ganeshie8 (ganeshie8):

oh denominators we dont subtract. below next steps for fraction part, see if it helps :

ganeshie8 (ganeshie8):

\(\huge \frac {29}{30} - \frac{26}{30} \) \(\huge \frac {1}{30}(29-26) \) \(\huge \frac {1}{30}(3) \) \(\huge \frac {3}{30} \) \(\huge \frac {1}{10} \)

ganeshie8 (ganeshie8):

once we have the common denominator of 30, we factored out 1/30 and subtracted 29-26

OpenStudy (anonymous):

okay, thank you!

ganeshie8 (ganeshie8):

np :D